# The consistency strength of determinacy when all sets are universally Baire

### v2

@@ 6.2. Comparison against a hull
Recall that $F$ is given by operators by Lemmas 5.16 and 5.22. We will use this fact together with ideas from the argument for condensation of the operators in [13, Lemma 2.27] to show that $E$ is given by operators as well. Let $\sigma\colon\operatorname{Ult}(\bar{\mathcal{M}},E)\rightarrow\operatorname{Ult}(\mathcal{M}_{\zeta},F)$ be the canonical factor map as defined above and let $\nu<\kappa_{0}$ be the least Woodin cardinal above $\delta_{0}$ in $\mathcal{M}_{\zeta}$ and $\bar{\mathcal{M}}$. Note that $Lp^{\infty}_{\omega}(\bar{\mathcal{M}}|\nu)=Lp^{\infty}_{\omega}(\mathcal{M}_{\zeta}|\nu)$ is suitable at $\kappa_{0}$ and $(\operatorname{Ord},\operatorname{Ord})$-iterable via the canonical iteration strategy obtained from $\Sigma$ by lifting the iteration trees to the background model $\mathcal{W}$ using that all extenders in $Lp^{\infty}_{\omega}(\mathcal{M}_{\zeta}|\nu)$ are fully backgrounded. Via this iteration strategy we can make ¯$\bar{\mathcal{M}}|\bar{\xi}$ and $\mathcal{M}_{\zeta}|\xi$ for $\xi=i(\bar{\xi})$ generic over iterates $\bar{\mathcal{N}}$ and $\mathcal{N}$ of $Lp^{\infty}_{\omega}(\bar{\mathcal{M}}|\nu)=Lp^{\infty}_{\omega}(\mathcal{M}_{\zeta}|\nu)$. Note that $\mathcal{M}_{\zeta}|\xi\in\operatorname{Ult}(\mathcal{M}_{\zeta},F)$ by coherence of $F$ and ¯$\bar{\mathcal{M}}|\bar{\xi}\in\operatorname{Ult}(\bar{\mathcal{M}},E)$ by Subclaim 3. So we in fact have $\bar{\mathcal{N}}\in\operatorname{Ult}(\bar{\mathcal{M}},E)$ and $\mathcal{N}\in\operatorname{Ult}(\mathcal{M}_{\zeta},F)$. Let $\bar{\pi}$ and $\pi$ denote the corresponding iteration embeddings, cf. Figure 9.
@@
Using the ideas from the proof of [13, Lemma 2.24], $S_{i}(\bar{\mathcal{N}})\in\operatorname{Ult}(\bar{\mathcal{M}},E)$ and analogously $S_{i}(\mathcal{N})\in\operatorname{Ult}(\mathcal{M}_{\zeta},F)$. In particular, $\sigma$ is defined on $S_{i}(\bar{\mathcal{N}})$. As $\bar{\pi}$ and $\pi$ are via the canonical iteration strategy obtained from $\Sigma$ by lifting trees to the background model, we have the same for their extensions to $S_{i}(Lp^{\infty}_{\omega}(\bar{\mathcal{M}}|\nu))$. By [13, Lemma 2.23],
@@
Finally, this implies that $E$ is the extender derived from $l$ and hence given by operators by the following computation. Let $a\in[\bar{\eta}]^{{<}\omega}$ and $A\subseteq[\kappa_{i}]^{|a|}$. Then, as the extender $F$ is given by operators and hence equal to the extender derived from the embedding $m$,
@@
As $E$ is given by operators, it agrees with the extender derived from the hull embedding $k\colon S_{i}(\bar{\mathcal{M}}|\kappa_{i})\rightarrow S_{i}(\bar{\mathcal{M}}|\bar{\xi})$. This implies ¯$\bar{\mathcal{M}}|(\bar{\xi}^{+})^{\operatorname{Ult}(\bar{\mathcal{M}},E)}\unlhd j_{E}(\bar{\mathcal{M}}|(\kappa_{i}^{+})^{\mathcal{M}})$ and hence finishes the argument that $E$ coheres ¯$\bar{\mathcal{M}}|\bar{\xi}$.
To show that $E$ is generically countably complete, let $G$ be generic over $\mathcal{W}$ for a forcing of size ${<}\kappa_{i}$ and let $\tau_{F}\colon\bigcup_{k}a_{k}\rightarrow\kappa_{i}$ be a function witnessing that $F$ is countably complete in $\mathcal{W}[G]$. Then $\tau_{E}$ given by $\tau_{E}(a)=x$ if and only if $\tau_{F}(i(a))=x$ witnesses that $E$ is countably complete in $\mathcal{W}[G]$ by the following argument: Let $\langle(a_{k},A_{k})\mid k<\omega\rangle$ be a sequence such that $(a_{k},A_{k})\in E$ for all $k<\omega$. Then $(i(a_{k}),A_{k})\in F$ and therefore $\tau_{F}\,"(i\,"a_{k})=\tau_{F}\,"i(a_{k})\in A_{k}$. Hence, $\tau_{E}\,"a_{k}\in A_{k}$, as desired.
As $E$ is given by operators and $\operatorname{Ult}(\mathcal{M}_{\zeta},E)|\bar{\xi}=\operatorname{Ult}(\bar{\mathcal{M}},E)|\bar{\xi}=\bar{\mathcal{M}}|\bar{\xi}$ by the previous subclaims, Lemma 5.23 yields $E\in M_{1}^{\#,\Sigma_{i}}(\bar{\mathcal{M}}|\bar{\xi})$. In particular, $E\in\bar{\mathcal{W}}$.
Finally, a straightforward search tree argument shows that $E$ is also generically countably complete in $\bar{\mathcal{W}}$ by absoluteness of well-foundedness. So $E$ must be added to $\bar{\mathcal{M}}$, the result of an $\mathcal{M}$-construction in $\bar{\mathcal{W}}$. But then we can argue that $E_{\alpha}^{\mathcal{M}^{*}}\neq\emptyset$.
@@
On both sides of the iteration, no extender with critical point ${\leq}\kappa_{i}$ is used on the main branch.
@@
By Claim 1 it suffices to show this for $\bar{\mathcal{M}}$. Suppose the $\bar{\mathcal{M}}$-side of the coiteration uses an extender with critical point ${\leq}\kappa_{i}$ and let $F$ be the first such extender that is used. Let $\mathcal{M}^{*}$ be the current stage of the iteration on the $\bar{\mathcal{M}}$-side, i.e., $F\in\mathcal{M}^{*}$, and let $\mathcal{U}$ be the iteration tree on $\bar{\mathcal{M}}$ of length $\lambda+1$ such that $\mathcal{M}^{*}=\mathcal{M}_{\lambda}^{\mathcal{U}}$. As $\mathcal{M}|\kappa_{i}=\bar{\mathcal{M}}|\kappa_{i}$, $F$ is indexed above $\kappa_{i}$. As in $\mathcal{M}$ no fully backgrounded extender overlaps one of the $\kappa_{j}$ for $j<\omega$, $F$ has critical point $\kappa_{j}$ for some $j\leq i$. We can argue as in the proof of Claim 1, cf. Case 2, to obtain that $F$ is certified.
@@
Recall that we suppose that $\kappa_{j}$ is fully strong in $\mathcal{M}$ and argue as in the case $\kappa_{j}<\kappa_{i}$ in the proof of the previous claim. So there is some extender $E_{\beta}^{\mathcal{M}}$ on the sequence of $\mathcal{M}$ with critical point $\kappa_{j}$ and $\beta$ above the index of $F$ (in the $\mathcal{M}^{*}$-sequence). Let $\beta$ be minimal with this property and write $\alpha$ for the $\mathcal{M}^{*}$-index of $F$. Then $\alpha$ is a cutpoint in $\operatorname{Ult}(\mathcal{M},E_{\beta}^{\mathcal{M}})$ and the $(\kappa_{j},\alpha)$-extender $E$ derived from the ultrapower embedding $\pi^{\mathcal{M}}_{E_{\beta}^{\mathcal{M}}}$ is on the $\mathcal{M}$-sequence by the initial segment condition. So $E_{\alpha}^{\mathcal{M}}=E$ and $E$ is on the sequence of $\mathcal{M}_{\nu}$ for some $\nu\geq\zeta$. Therefore, even if $F$ is used in the comparison against $\mathcal{M}_{\zeta}$ it will not be used on the main branch in the comparison of $\bar{\mathcal{M}}$ against the construction of $\mathcal{M}$ as this part of the coiteration gets collapsed when moving on to the comparison with $\mathcal{M}_{\nu}$.
@@ 6.3. The construction converges and the $\kappa_{i}$ are strong all the way
$\mathcal{M}$ is a proper class premouse, i.e., the $\mathcal{M}$-construction in $\mathcal{W}$ does not break down.
@@
Let $(\mathcal{M}_{\xi},\mathcal{N}_{\xi}\mid\xi<\zeta),\mathcal{N}_{\zeta}$ be the models in the construction of $\mathcal{M}$ and suppose toward a contradiction that the $\mathcal{M}$-construction breaks down at stage $\zeta$. That means $\mathcal{M}_{\xi}=\mathcal{C}_{\omega}(\mathcal{N}_{\xi})$ for $\xi<\zeta$, $\zeta$ is a successor ordinal and $\mathcal{C}_{\omega}(\mathcal{N}_{\zeta})$ is not defined. Say $\mathcal{N}_{\zeta}=(J_{\alpha}^{\vec{E}},\in,\vec{E},F)$ for some generically countably complete extender $F$ with critical point $\kappa_{i}$ for some $i<\omega$ such that $\delta_{i+1}<\alpha$. We may assume that all $\kappa_{j}$ for $j<i$ are strong in $\mathcal{N}_{\zeta}$ as we can otherwise run the proof of Lemma 6.6 below. Let $\bar{\mathcal{W}}$ be the transitive collapse of a hull of $\mathcal{W}|\Omega$ for some sufficiently large $\Omega$ such that $\mathcal{W}|\kappa_{i}\subseteq\bar{\mathcal{W}}$ and $|\bar{\mathcal{W}}|=\kappa_{i}$. Let
$\pi\colon\bar{\mathcal{W}}\rightarrow\mathcal{W}$
be the uncollapse embedding, $\bar{\mathcal{M}}$ the collapse of $\mathcal{M}|\Omega$, and $(\bar{\kappa}_{j},\bar{\delta}_{j}\mid j<\omega)$ the preimages of $(\kappa_{j},\delta_{j}\mid j<\omega)$ (so in particular, $\bar{\kappa}_{j}=\kappa_{j}$ and $\bar{\delta}_{j}=\delta_{j}$ for $j\leq i$). By Lemma 6.4 there is an iterate $\mathcal{M}^{*}$ of $\bar{\mathcal{M}}$ such that the main branch from $\bar{\mathcal{M}}$ to $\mathcal{M}^{*}$ does not drop and if we let
@@
Let $(\delta_{i}\mid i<\omega)$ be the sequence of Woodin cardinals in $\mathcal{W}$ and for each $i<\omega$, let $\kappa_{i}$ be the least ${<}\delta_{i+1}$-strong cardinal above $\delta_{i}$. Then $\kappa_{i}$ is a strong cardinal in $\mathcal{M}$ for all $i<\omega$.
@@
Suppose there is an $i<\omega$ such that $\kappa_{i}$ is not $\gamma$-strong for some $\gamma$. We may assume that $\gamma$ is a cardinal. Let $i$ be minimal with this property, i.e., suppose that all cardinals $\kappa_{j}$ with $j<i$ are fully strong in $\mathcal{M}$. Let $\bar{\mathcal{W}}$ be the transitive collapse of a hull of $\mathcal{W}|\Omega$ for some sufficiently large $\Omega$ such that $\mathcal{W}|\kappa_{i}\subseteq\bar{\mathcal{W}}$ and $|\bar{\mathcal{W}}|=\kappa_{i}$. Let
$\pi\colon\bar{\mathcal{W}}\rightarrow\mathcal{W}$
be the uncollapse embedding, $\bar{\mathcal{M}}$ the collapse of $\mathcal{M}|\Omega$, and $(\bar{\kappa}_{j},\bar{\delta}_{j},\bar{\gamma}\mid j<\omega)$ the preimages of $(\kappa_{j},\delta_{j},\gamma\mid j<\omega)$ (so in particular, $\bar{\kappa}_{j}=\kappa_{j}$ and $\bar{\delta}_{j}=\delta_{j}$ for $j\leq i$). By Lemma 6.4 there is an iterate $\mathcal{M}^{*}$ of $\bar{\mathcal{M}}$ such that the main branch from $\bar{\mathcal{M}}$ to $\mathcal{M}^{*}$ does not drop and if we let
@@
denote the corresponding iteration embedding, then $\mathcal{M}^{*}=\mathcal{M}_{\chi}$ for some model $\mathcal{M}_{\chi}$ in the construction of $\mathcal{M}$ and the critical point of $i$ is above $\kappa_{i}$. As in the proof of Claim 1 in the proof of Lemma 6.4 we have $\mathcal{M}^{*}\cap\operatorname{Ord}<\delta_{i+1}$. Recall that $\kappa_{i}$ is ${<}\delta_{i+1}$-strong in $\mathcal{M}$, so there are extenders on the $\mathcal{M}$-sequence with arbitrarily large length $\eta$ between $i(\bar{\gamma})$ and $\delta_{i+1}$. Let $F$ be an extender on the $\mathcal{M}_{\chi}$-sequence with critical point $\kappa_{i}$ and length $\eta$ between $i(\bar{\gamma})$ and $\delta_{i+1}$. Now consider the extender $E$ with critical point $\kappa_{i}$ given by
$(a,A)\in E\text{ iff }(i(a),A)\in F$
for $a\in[\operatorname{Ord}]^{<\omega}$ and $A\subseteq[\kappa_{i}]^{|a|}$. As the critical point of $i$ is above $\kappa_{i}$, we have $\mathcal{M}\cap\mathcal{P}(\kappa_{i})=\bar{\mathcal{M}}\cap\mathcal{P}(\kappa_{i})$ and this defines an extender $E$ over $\bar{\mathcal{M}}$. Let $\bar{\xi}$ be the least $i$-weak cutpoint above $\kappa_{i}$ in $\operatorname{Ult}(\bar{\mathcal{M}},E)$. Recall that an ordinal $\alpha$ is an $i$-weak cutpoint above $\kappa_{i}$ if all extenders overlapping $\alpha$ have critical point $\kappa_{j}$ for some $j<i$. So $\bar{\eta}=(\bar{\xi}^{+})^{\operatorname{Ult}(\bar{\mathcal{M}},E)}$ is the ordinal where $E$ would be indexed if it is on the sequence of $\bar{\mathcal{M}}$. As in the proof of Claim 1 in the proof of Lemma 6.4 we obtain the following claim.
@@
$E$ coheres ¯$\bar{\mathcal{M}}|\bar{\eta}$ and is generically countably complete in $\bar{\mathcal{W}}$.
Hence, $E$ must be added to $\bar{\mathcal{M}}$, the result of an $\mathcal{M}$-construction in $\bar{\mathcal{W}}$, contradicting the fact that $\kappa_{i}$ is not $\bar{\gamma}$-strong in $\bar{\mathcal{M}}$. ∎
As all $\delta_{i}$, $i<\omega$, remain Woodin in $\mathcal{M}$, Lemma 6.6 yields that $\delta_{\omega}=\sup_{i<\omega}\delta_{i}$ is a limit of Woodin cardinals and a limit of strong cardinals in $\mathcal{M}$. This finishes the proof of Theorem 1.1.

### v3

@@ 6.2. Comparison against a hull
Recall that $E$ is given by operators by Lemmas 5.16 and 5.22. We will use this fact together with ideas from the argument for condensation of the operators in [14, Lemma 2.27] to show that $F$ is given by operators as well. Let $\sigma\colon\operatorname{Ult}(\bar{\mathcal{M}},F)\rightarrow\operatorname{Ult}(\mathcal{M}_{\zeta},E)$ be the canonical factor map as defined above and let $\nu<\kappa_{0}$ be the least Woodin cardinal above $\delta_{0}$ in $\mathcal{M}_{\zeta}$ and $\bar{\mathcal{M}}$. Note that $Lp^{\infty}_{\omega}(\bar{\mathcal{M}}|\nu)=Lp^{\infty}_{\omega}(\mathcal{M}_{\zeta}|\nu)$ is suitable at $\kappa_{0}$ and $(\operatorname{Ord},\operatorname{Ord})$-iterable via the canonical iteration strategy obtained from $\Sigma$ by lifting the iteration trees to the background model $\mathcal{W}$ using that all extenders in $Lp^{\infty}_{\omega}(\mathcal{M}_{\zeta}|\nu)$ are fully backgrounded. Via this iteration strategy we can make ¯$\bar{\mathcal{M}}|\bar{\xi}$ and $\mathcal{M}_{\zeta}|\xi$ for $\xi=i(\bar{\xi})$ generic over iterates $\bar{\mathcal{N}}$ and $\mathcal{N}$ of $Lp^{\infty}_{\omega}(\bar{\mathcal{M}}|\nu)=Lp^{\infty}_{\omega}(\mathcal{M}_{\zeta}|\nu)$. Note that $\mathcal{M}_{\zeta}|\xi\in\operatorname{Ult}(\mathcal{M}_{\zeta},E)$ by coherence of $E$ and ¯$\bar{\mathcal{M}}|\bar{\xi}\in\operatorname{Ult}(\bar{\mathcal{M}},F)$ by Subclaim 3. So we in fact have $\bar{\mathcal{N}}\in\operatorname{Ult}(\bar{\mathcal{M}},F)$ and $\mathcal{N}\in\operatorname{Ult}(\mathcal{M}_{\zeta},E)$. Let $\bar{\pi}$ and $\pi$ denote the corresponding iteration embeddings, cf. Figure 9.
@@
Using the ideas from the proof of [14, Lemma 2.24], $S_{i}(\bar{\mathcal{N}})\in\operatorname{Ult}(\bar{\mathcal{M}},F)$ and analogously $S_{i}(\mathcal{N})\in\operatorname{Ult}(\mathcal{M}_{\zeta},E)$. In particular, $\sigma$ is defined on $S_{i}(\bar{\mathcal{N}})$. As $\bar{\pi}$ and $\pi$ are via the canonical iteration strategy obtained from $\Sigma$ by lifting trees to the background model, we have the same for their extensions to $S_{i}(Lp^{\infty}_{\omega}(\bar{\mathcal{M}}|\nu))$. By [14, Lemma 2.23],
@@
Finally, this implies that $F$ is the extender derived from $l$ and hence given by operators by the following computation. Let $a\in[\bar{\eta}]^{{<}\omega}$ and $A\subseteq[\kappa_{i}]^{|a|}$. Then, as the extender $E$ is given by operators and hence equal to the extender derived from the embedding $m$,
@@
As $F$ is given by operators, it agrees with the extender derived from the hull embedding $k\colon S_{i}(\bar{\mathcal{M}}|\kappa_{i})\rightarrow S_{i}(\bar{\mathcal{M}}|\bar{\xi})$. This implies ¯$\bar{\mathcal{M}}|(\bar{\xi}^{+})^{\operatorname{Ult}(\bar{\mathcal{M}},F)}\unlhd j_{F}(\bar{\mathcal{M}}|(\kappa_{i}^{+})^{\mathcal{M}})$ and hence finishes the argument that $F$ coheres ¯$\bar{\mathcal{M}}|\bar{\xi}$.
To show that $F$ is generically countably complete, let $G$ be generic over $\mathcal{W}$ for a forcing of size ${<}\kappa_{i}$ and let $\tau_{E}\colon\bigcup_{k}a_{k}\rightarrow\kappa_{i}$ be a function witnessing that $E$ is countably complete in $\mathcal{W}[G]$. Then $\tau_{F}$ given by $\tau_{F}(a)=x$ if and only if $\tau_{E}(i(a))=x$ witnesses that $F$ is countably complete in $\mathcal{W}[G]$ by the following argument: Let $\langle(a_{k},A_{k})\mid k<\omega\rangle$ be a sequence such that $(a_{k},A_{k})\in F$ for all $k<\omega$. Then $(i(a_{k}),A_{k})\in E$ and therefore $\tau_{E}\,"(i\,"a_{k})=\tau_{E}\,"i(a_{k})\in A_{k}$. Hence, $\tau_{F}\,"a_{k}\in A_{k}$, as desired.
As $F$ is given by operators and $\operatorname{Ult}(\mathcal{M}_{\zeta},F)|\bar{\xi}=\operatorname{Ult}(\bar{\mathcal{M}},F)|\bar{\xi}=\bar{\mathcal{M}}|\bar{\xi}$ by the previous subclaims, Lemma 5.23 yields $F\in M_{1}^{\#,\Sigma_{i}}(\bar{\mathcal{M}}|\bar{\xi})$. In particular, $F\in\bar{\mathcal{W}}$.
Finally, a straightforward search tree argument shows that $F$ is also generically countably complete in $\bar{\mathcal{W}}$ by absoluteness of well-foundedness. So $F$ must be added to $\bar{\mathcal{M}}$, the result of an $\mathcal{M}$-construction in $\bar{\mathcal{W}}$. But then we can argue that $E_{\alpha}^{\mathcal{M}^{*}}\neq\emptyset$.
@@
The following claim is another key observation in the comparison argument.
@@
On both sides of the iteration, no extender with critical point ${\leq}\kappa_{i}$ is used.
@@
By Claim 1 it suffices to show this for $\bar{\mathcal{M}}$. Suppose the $\bar{\mathcal{M}}$-side of the coiteration uses an extender with critical point ${\leq}\kappa_{i}$ and let $F$ be the first such extender that is used. Let $\mathcal{M}^{*}$ be the current stage of the iteration on the $\bar{\mathcal{M}}$-side, i.e., $F\in\mathcal{M}^{*}$, and let $\mathcal{U}$ be the iteration tree of length $\lambda+1$ on some model ¯$\bar{\mathcal{M}}_{\xi}$ in the construction of $\bar{\mathcal{M}}$ such that $\mathcal{M}^{*}=\mathcal{M}_{\lambda}^{\mathcal{U}}$. As $\mathcal{M}|\kappa_{i}=\bar{\mathcal{M}}|\kappa_{i}=\bar{\mathcal{M}}_{\xi}|\kappa_{i}$, $F$ is indexed above $\kappa_{i}$. As in $\mathcal{M}$ no fully backgrounded extender overlaps $\kappa_{j}$ for any $j<\omega$, $F$ has critical point $\kappa_{j}$ for some $j\leq i$. We can argue as in the proof of Claim 1, cf. Case 2, to obtain that $F$ is certified.
@@
Recall that we suppose that $\kappa_{j}$ is fully strong in $\mathcal{M}$ and argue as in the case $\kappa_{j}<\kappa_{i}$ in the proof of the previous claim. So there is some extender $E_{\beta}^{\mathcal{M}}$ on the sequence of $\mathcal{M}$ with critical point $\kappa_{j}$ and $\beta$ above the index of $F$ (in the $\mathcal{M}^{*}$-sequence). Let $\beta$ be minimal with this property and write $\alpha$ for the $\mathcal{M}^{*}$-index of $F$. Then $\alpha$ is a cutpoint in $\operatorname{Ult}(\mathcal{M},E_{\beta}^{\mathcal{M}})$ and the $(\kappa_{j},\alpha)$-extender $E$ derived from the ultrapower embedding $\pi^{\mathcal{M}}_{E_{\beta}^{\mathcal{M}}}$ is on the $\mathcal{M}$-sequence by the initial segment condition. So $E_{\alpha}^{\mathcal{M}}=E$ and $E$ is on the sequence of $\mathcal{M}_{\nu}$ for some $\nu\geq\zeta$. Therefore, even if $F$ is used in the comparison of the constructions of $\bar{\mathcal{M}}$ and $\mathcal{M}$ it will not be used in the iteration from ¯$\bar{\mathcal{M}}_{\xi}$ to $\mathcal{M}^{*}$ as the part of the coiteration using $F$ gets collapsed when moving on to the comparison with $\mathcal{M}_{\nu}$.
@@ 6.3. The construction converges and the $\kappa_{i}$ are strong all the way
Let $i<\omega$ and suppose $(\mathsf{IH})_{i}$. Then $\mathcal{M}^{(i+1)}$ is a proper class premouse, i.e., the $\mathcal{M}$-construction in $\mathcal{W}$ does not break down when restricting Clause (2c) to extender with critical points in $\{\kappa_{j}\mid j\leq i\}$.
@@
Let $(\mathcal{M}_{\xi},\mathcal{N}_{\xi}\mid\xi<\zeta),\mathcal{N}_{\zeta}$ be the models in the construction of $\mathcal{M}^{(i+1)}$ and suppose toward a contradiction that the $\mathcal{M}$-construction breaks down at stage $\zeta$. That means $\mathcal{M}_{\xi}=\mathcal{C}_{\omega}(\mathcal{N}_{\xi})$ for $\xi<\zeta$, $\zeta$ is a successor ordinal and $\mathcal{C}_{\omega}(\mathcal{N}_{\zeta})$ is not defined. Say $\mathcal{M}^{(i+1)}=\mathcal{N}_{\zeta}=(J_{\alpha}^{\vec{E}},\in,\vec{E},F)$ for some generically countably complete extender $F$ with critical point $\kappa_{i}$ for some $i<\omega$ such that $\delta_{i+1}<\alpha$. Let
$\pi\colon\bar{\mathcal{W}}\rightarrow\mathcal{W}|\Omega$
for some sufficiently large $\Omega>\mathcal{M}^{(i+1)}\cap\operatorname{Ord}$ be such that letting $\operatorname{crit}(\pi)=\nu$, $\mathcal{W}|\kappa_{i}\subseteq\bar{\mathcal{W}}$, $(\kappa_{j},\delta_{j}\mid j<\omega)\in\operatorname{rng}(\pi)$, the sequence of models of the construction of $\mathcal{M}^{(i+1)}$ is in the range of $\pi$, $\nu\in(\kappa_{i},\delta_{i+1})$ is an inaccessible cardinal in $\mathcal{W}$, and $\pi(\nu)=\delta_{i+1}$. Let $\bar{\mathcal{M}}$ be the collapse of $\mathcal{M}^{(i+1)}$ and let $(\bar{\kappa}_{j},\bar{\delta}_{j}\mid j<\omega)$ be the preimages of $(\kappa_{j},\delta_{j}\mid j<\omega)$ (so in particular, $\bar{\kappa}_{j}=\kappa_{j}$ and $\bar{\delta}_{j}=\delta_{j}$ for $j\leq i$). By Lemma 6.4 there is an iterate $\mathcal{M}^{*}$ of $\bar{\mathcal{M}}$ such that the main branch from $\bar{\mathcal{M}}$ to $\mathcal{M}^{*}$ does not drop and if we let
@@
The next lemma proves $(\mathsf{IH}.3)_{i+1}$ from $(\mathsf{IH})_{i}$, finishing the inductive argument for $(\mathsf{IH})_{i+1}$ from $(\mathsf{IH})_{i}$.
@@
Let $(\delta_{i}\mid i<\omega)$ be the sequence of Woodin cardinals in $\mathcal{W}$ and for each $i<\omega$, let $\kappa_{i}$ be the least ${<}\delta_{i+1}$-strong cardinal above $\delta_{i}$. Let $i<\omega$ and assume $(\mathsf{IH})_{i}$. Then $\kappa_{j}$ is a strong cardinal in $\mathcal{M}^{(i+1)}$ for all $j\leq i$.
@@
Suppose there is a $j\leq i$ such that $\kappa_{j}$ is not $\gamma$-strong in $\mathcal{M}^{(i+1)}$ for some $\gamma$. We may assume that $\gamma$ is a cardinal. Let $j$ be minimal with this property, i.e., suppose that all cardinals $\kappa_{l}$ with $l<j$ are fully strong in $\mathcal{M}^{(i+1)}$. By Lemma 6.5 $\mathcal{M}^{(i+1)}$ is a proper class model. We may assume $j=i$ as the argument for $j<i$ is similar. Let
$\pi\colon\bar{\mathcal{W}}\rightarrow\mathcal{W}|\Omega$
for some sufficiently large $\Omega$ be such that letting $\operatorname{crit}(\pi)=\nu$, $\mathcal{W}|\kappa_{i}\subseteq\bar{\mathcal{W}}$, $(\kappa_{j},\delta_{j}\mid j<\omega)\in\operatorname{rng}(\pi)$, the sequence of models of the construction of $\mathcal{M}^{(i+1)}|\Omega$ is in the range of $\pi$, $\nu\in(\kappa_{i},\delta_{i+1})$ is an inaccessible cardinal in $\mathcal{W}$, and $\pi(\nu)=\delta_{i+1}$. Let $\bar{\mathcal{M}}$ be the collapse of $\mathcal{M}^{(i+1)}|\Omega$ and $(\bar{\kappa}_{j},\bar{\delta}_{j},\bar{\gamma}\mid j<\omega)$ the preimages of $(\kappa_{j},\delta_{j},\gamma\mid j<\omega)$ (so in particular, $\bar{\kappa}_{j}=\kappa_{j}$ and $\bar{\delta}_{j}=\delta_{j}$ for $j\leq i$). By Lemma 6.4 there is an iterate $\mathcal{M}^{*}$ of $\bar{\mathcal{M}}$ such that the main branch from $\bar{\mathcal{M}}$ to $\mathcal{M}^{*}$ does not drop and if we let
@@
denote the corresponding iteration embedding, then $\mathcal{M}^{*}=\mathcal{M}_{\chi}$ for some model $\mathcal{M}_{\chi}$ in the construction of $\mathcal{M}^{(i+1)}$ and the critical point of $i$ is above $\kappa_{i}$. As in the proof of Claim 1 in the proof of Lemma 6.4 we have $\mathcal{M}^{*}\cap\operatorname{Ord}<\delta_{i+1}$. Recall that $\kappa_{i}$ is ${<}\delta_{i+1}$-strong in $\mathcal{M}^{(i+1)}$, so there are extenders on the $\mathcal{M}^{(i+1)}$-sequence with arbitrarily large length $\eta$ between $i(\bar{\gamma})$ and $\delta_{i+1}$. Let $E$ be an extender on the $\mathcal{M}_{\chi}$-sequence with critical point $\kappa_{i}$ and length $\eta$ between $i(\bar{\gamma})$ and $\delta_{i+1}$. Now consider the extender $F$ with critical point $\kappa_{i}$ given by
$(a,A)\in F\text{ iff }(i(a),A)\in E$
for $a\in[\operatorname{Ord}]^{<\omega}$ and $A\subseteq[\kappa_{i}]^{|a|}$. As the critical point of $i$ is above $\kappa_{i}$, we have $\mathcal{M}\cap\mathcal{P}(\kappa_{i})=\bar{\mathcal{M}}\cap\mathcal{P}(\kappa_{i})$ and this defines an extender $F$ over $\bar{\mathcal{M}}$. Let $\bar{\xi}$ be the least $i$-weak cutpoint above $\kappa_{i}$ in $\operatorname{Ult}(\bar{\mathcal{M}},F)$. Recall that an ordinal $\alpha$ is an $i$-weak cutpoint above $\kappa_{i}$ if all extenders overlapping $\alpha$ have critical point $\kappa_{j}$ for some $j<i$. So $\bar{\eta}=(\bar{\xi}^{+})^{\operatorname{Ult}(\bar{\mathcal{M}},F)}$ is the ordinal where $F$ would be indexed if it is on the sequence of $\bar{\mathcal{M}}$. As in the proof of Claim 1 in the proof of Lemma 6.4 we obtain the following claim.
@@
$F$ coheres ¯$\bar{\mathcal{M}}|\bar{\eta}$ and is generically countably complete in $\bar{\mathcal{W}}$.
Hence, $F$ must be added to $\bar{\mathcal{M}}$, the result of an $\mathcal{M}^{(i+1)}$-construction in $\bar{\mathcal{W}}$, contradicting the fact that $\kappa_{i}$ is not $\bar{\gamma}$-strong in $\bar{\mathcal{M}}$. ∎
As all $\delta_{i}$, $i<\omega$, remain Woodin in $\mathcal{M}$, Lemmas 6.5 and 6.6 inductively yield that $\delta_{\omega}=\sup_{i<\omega}\delta_{i}$ is a limit of Woodin cardinals and a limit of strong cardinals in the proper class model $\mathcal{M}$. This finishes the proof of Theorem 1.1.